OpenStudy (anonymous):

A Ferris Wheel 22.0m in diameter rotates once every 12.5s. What is the ratio of a persons apperenet weight to her real weight (a.) at the top, and (b) at the bottom?

OpenStudy (loser66):

@Jhannybean @ybarrap

OpenStudy (jhannybean):

@bahrom7893

OpenStudy (bahrom7893):

Lag.. anyway..

OpenStudy (bahrom7893):

Apparent weight: N = mg - F_up

OpenStudy (bahrom7893):

So basically, you have two forces acting on the body - the force of gravity pulling it down, and the seat pushing it up.

OpenStudy (bahrom7893):

The net force, F = m*a, and we're going to have to use centripetal acceleration in this problem, since it's a Ferris wheel (sorry just writing down random notes for myself, I'm pretty bad at physics and I generally like to gather as much as I can before even trying to solve something)

OpenStudy (bahrom7893):

So, centripetal acceleration is: a = w^2 * r (w is omega in this case), where w = 2pi/t

OpenStudy (jhannybean):

Ah, thought this belonged in the physics section. Hmm.

OpenStudy (bahrom7893):

Ok.. now I'm kinda lost... Aren't you in the physics section?????

OpenStudy (jhannybean):

2* initial momentum/time?

OpenStudy (bahrom7893):

Okay, I just did a bit of googling... andd found this problem.. enjoy (lol sorry too lazy to try to follow it): http://answers.yahoo.com/question/index?qid=20100427151843AAd3t7s